OpenStudy (anonymous):

Both 2(cos10+isin10) and 2(cos100+isin100) are nth roots of a particular number. What is the smallest possible value of n?

OpenStudy (baru):

4 ? 100+100+100+100=360+40 10+10+10+10=40